### v2

# Stars and quark stars in bumblebee gravity

Abstract
In this work, the interior spacetime of stars is built in a Lorentz symmetry breaking model called bumblebee gravity. Firstly, we calculated the modified Tolman-Oppenheimer-Volkoff equation in this context of modified gravity. Then we show that the bumblebee field, responsible for the symmetry breaking, increases the star mass-radius relation when it assumes its vacuum expectation value. When compared to the general relativity mass-radius relation, a Lorentz symmetry breaking context, like the bumblebee gravity, could provide more massive stars, surpassing the $2.5M_{\odot}$ limit as the interior of the star is described by quark matter with the MIT bag model.
Neutron Star, Quark Gluon Plasma, Lorentz Symmetry Breaking
I Introduction
[…]
However, when gravity is at stake, one should take into account a mechanism of spontaneous symmetry breaking Bluhm2005 ; Bluhm2008 in order to maintain the geometric constraints and conservation laws from general relativity. In this sense, when gravity is not weak, bumblebee models provide the Lorentz symmetry breaking (spontaneously broken) as the bumblebee field (a vector or a tensor field) reaches its nonzero vacuum expectation value (VEV). In the gravitational sector, bumblebee models are studied in curved spacetimes Kostelecky2004 ; Bailey2006 ; Filho:2022yrk , gravitational waves Maluf:2014dpa ; Kostelecky:2016kkn ; Kostelecky:2016kfm ; Amarilo:2023wpn , black hole physics Bertolami ; Casana ; Euclides2020 ; Maluf_Neves , wormholes Ovgun2020 ; Rondineli2019 , and cosmology Capelo:2015ipa ; ONeal ; Petrov ; Neves:2022qyb ; Maluf:2021eyu ; Maluf:2021lwh .
In this work, we computed the interior geometry of a star or quark star in the bumblebee context. Quark stars are important because, in conditions of extreme energy density, the strong nuclear force results in matter composed of quarks and gluons, as evidenced by relativistic heavy-ion collisions Gardim:2019xjs . Recent data suggests that observations of compact stars correspond with the description of quark stars, or at least, are consistent with quark cores within neutron stars Annala:2019puf . For that purpose, we assume that the Casana et al. Casana solution, also called Schwarzschild-like black hole, describes the exterior of a static and noncharged star in bumblebee gravity. Like the general relativity case, in which the interior solution matches the Schwarzschild metric at the radius of the star, in the case studied here we adopt the same strategy to build the interior solution. The modified Tolman-Oppenheimer-Volkoff (TOV) equation is calculated. Whether in general relativity or in the bumblebee model, the TOV equation provides the hydrostatic equilibrium description of fluids inside compact objects like neutron stars or quark stars. Two types of fluid will be used to describe the matter inside the star and to solve the modified TOV. In the first scenario, we employ a fluid with a constant density (incompressible) as a didactic case. The second one is the quark matter described by the MIT bag model, something more realistic than the first case.
The MIT bag model Chodos:1974je is a simplified way to describe how quarks and gluons are confined in protons and neutrons. It is based on the idea that the energy of this state comes from the lowest energy state of the quarks and gluons, plus the vacuum of this state, which is different from the vacuum outside a small region. We refer to this energy as the bag constant, $\mathcal{B}$, representing the energy within the bag (a small region) that confines quarks and gluons. In turn, this same energy generates a confinement force, balancing the pressure of the quarks inside the bag with the pressure of the external vacuum. With this description, one can construct an equation of state in which there is a contribution of this vacuum energy. Although this model is a phenomenological description and does not account for the details of quantum chromodynamics interaction, it can aid in understanding qualitatively the behavior of matter under extreme density conditions, such as in compact stars like neutron stars and quark stars Itoh:1970uw ; Farhi:1984qu ; Witten:1984rs ; Fraga:2013qra .
In this sense, quark matter in astrophysics has been subject of many papers. In modified theories of gravity, we see works in Einstein-Gauss-Bonnet gravity Gammon:2023uss , $f(R,L_{m},T)$ gravity Tangphati:2024war , Weyl geometric gravity Haghani:2023nrm and many others (see Ref. Olmo:2019flu for a review). Quark stars are studied in modified gravity due to the possibility of having larger mass-radius relations in these theories (when compared to general relativity). Recent data of gravitational waves events, like GW 190814 LIGO , points out that one of the components of the coalescence between a black hole and a compact star presents mass $M=2.59^{+0.08}_{-0.09}M_{\odot}$, where $M_{\odot}$ is the solar mass. However, using the MIT bag model and reasonable values for the bag constant (as we will see), general relativity is not able to surpass the $2.4M_{\odot}$ value for a static and noncharged star. On the other hand, a star in the bumblebee gravity could have $M=2.53M_{\odot}$ or even larger values.
This article is structured as follows: In Section II the adopted bumblebee model is presented, the exterior solution, Schwarzschild-like spacetime, is briefly discussed. The analytic interior solution is built. Section III discusses two types of fluid for solving the modified TOV equation. In the incompressible case, approximate expressions for the pressure and the metric inside the star are shown. The second type of fluid under study is the quark matter that the MIT bag model describes. The mass-radius result is exhibited, and we can see the enhancement of this relation due to the bumblebee field. The final comments are in Section IV. In our calculations we adopt geometrized units, i.e., $G=c=1$, where $G$ is the Newtonian constant, and $c$ is speed of light in vacuum. As we present the results, the units in terms of MeV/fm3 are restored.
II Stars in bumblebee gravity
II.1 The adopted model
Bumblebee models are simple mechanisms for studying, in the gravitational sector, the spontaneous breaking of (arguably) the most important symmetry in physics, the Lorentz symmetry. As mentioned in Introduction, these models have a nontrivial VEV, which could affect the dynamics of other fields coupled to the bumblebee field. Also bumblebee models preserve both geometric structures and conservation laws compatible with general relativity Kostelecky2004 ; Bluhm2005 .
The model adopted here regards a vector field $B_{\mu}$, the bumblebee field (in a torsion-free spacetime), given by the following action:
$S_{B}$ $=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2\kappa}+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}$
$\ \ \ -V(B^{\mu}B_{\mu}\pm b^{2})+\mathcal{L}_{M}\bigg{]},$ (1)
where $\kappa=8\pi G/c^{4}$ is the gravitational coupling constant, and $\xi$ (with mass dimension $[\xi]=M^{-1}$) plays the role of a coupling constant that accounts for the nonminimum interaction between the bumblebee field and the geometry Bailey2006 . Another important ingredient in the action is the field strength, defined as
$B_{\mu\nu}\equiv\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}.$ (2)
The Lagrangian $\mathcal{L}_{M}$ describes the matter content, which will be treated as a perfect fluid in the next section. Lastly, the potential $V$ is responsible for triggering spontaneously the Lorentz violation as the bumblebee field assumes a nonzero VEV $\left\langle B_{\mu}\right\rangle\equiv b_{\mu}$, obeying the condition $B^{\mu}B_{\mu}=\mp b^{2}$. It is worth emphasizing that the constant $b^{2}$ is a real (thus positive) number, and the $\pm$ sign in Eq. (1) means that $b_{\mu}$ is timelike or spacelike, respectively, in agreement with the metric signature adopted in this article, namely (,+,+,+)(-,+,+,+)( - , + , + , + ).
The modified gravitational field equations in the bumblebee context are directly obtained by varying the action (1) with respect to the metric tensor $g_{\mu\nu}$ (as the bumblebee field $B_{\mu}$ is held fixed). This calculation leads to
$G_{\mu\nu}=$ $\ R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa\left(T^{B}_{\mu\nu}+T^{M}_{\mu\nu}\right)$
$=$ $\ \kappa\left[2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu}\right]$
$+\xi\bigg{[}\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}$
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
$-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\bigg{]}+\kappa T^{M}_{\mu\nu}.$ (3)
The tensor $G_{\mu\nu}$ is the usual Einstein tensor, and the operator means derivative with respect to the potential argument (consequently, throughout this article is also derivative with respect to the radial coordinate), and $\nabla_{\mu}$ is the covariant derivative. The bumblebee energy-momentum tensor and the matter energy-momentum tensor are $T^{B}_{\mu\nu}$ and $T^{M}_{\mu\nu}$, respectively. In order to solve Eq. (3) and to build a spacetime metric, we should choose a bumblebee potential $V$ and a metric Ansatz with the desired spacetime symmetry. Here, we study static stars, therefore we should adopt the the general spherically symmetric Ansatz. From these choices, a set of equations is obtained, and a full metric is written. In our case, the interior metric of a star.
The bumblebee model action (1) also gives us the equation of motion for the field $B_{\mu}$. That is, by varying (1) with respect to the field $B_{\mu}$, one has
$\nabla_{\mu}B^{\mu\nu}=J_{B}^{\nu}+J_{M}^{\nu},$ (4)
where $J_{M}^{\nu}$, due to the matter content, is some sort of source for the bumblebee field, and
$J_{B}^{\nu}=2\left(V^{\prime}B^{\nu}-\frac{\xi}{2\kappa}B_{\mu}R^{\mu\nu}\right)$ (5)
is the term that describes the self-interaction of the field $B_{\mu}$.
II.2 The exterior solution
In Casana et al. Casana , a vacuum and static solution in the bumblebee context was built from the spherically symmetric Ansatz in the $(t,r,\theta,\phi)$ coordinates, i.e.,
$g_{\mu\nu}=\mbox{diag}(-e^{2\alpha(r)},e^{2\beta(r)},r^{2},r^{2}\sin^{2}\theta).$ (6)
The authors also have chosen the following radial bumblebee field
$B_{\mu}=b_{\mu}=(0,b_{r}(r),0,0),$ (7)
where
$b_{r}(r)=\left|b\right|e^{\beta(r)}.$ (8)
From those choices, we see that the condition $b_{\mu}b^{\mu}=b^{2}=\mbox{constant}$ is fully satisfied. It is worth noting that the vacuum condition of the field (7) leads then to a bumblebee field frozen in its VEV, that is, $V=0$, and the assumption $V^{\prime}=0$ guarantees that the field keeps on the minimum of the potential. Another consequence of that particular field choice is that the field strength is zero ($B_{\mu\nu}=0$).
From both the Ansatz (6) and the field (7), according to Ref. Casana , the metric able to describe a static and neutral black hole in vacuum (or alternatively the exterior region of a nonrotating star) in the bumblebee context is
$ds^{2}=$ $-\left(1-\frac{2M}{r}\right)dt^{2}+(1+\ell)\left(1-\frac{2M}{r}\right)^{-1}dr^{2}$
$+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (9)
That is to say, the spacetime described by Eq. (9), also called Schwarzschild-like spacetime, is a spherically symmetric solution of the modified field equations (3) in vacuum, that is, $T_{\mu\nu}^{M}=0$. The parameter $M$ is conceived of as the black hole mass (or stellar mass in this article). The Lorentz symmetry is broken by the Lorentz-violating parameter $\ell=\xi b^{2}$, it is the presence of the bumblebee field in the geometry. Of course, for $\ell=0$, the Schwarzschild metric is restored in the general relativity context. Also, Casana et al., in the same work, evaluated the light deflection, time delay of light, and the perihelion advance of the planet Mercury in order to obtain upper bounds on the parameter $\ell$. Accordingly, the most stringent upper bound is $\ell<10^{-15}$. However, such a limit comes from Solar system experiments. In this article, due to high pressures and densities, we are going to relax this upper bound, allowing then higher values for the Lorentz-violating parameter.
Lastly, as pointed out in Casana , the bumblebee field acts on the geometry modifying measurable parameters. However, as the event horizon localization is given by $g^{rr}=e^{-2\beta}=0$, we can see that the event horizon radius $r_{+}=2M$ is the same in the Schwarzschild and Schwarzschild-like spacetimes.
II.3 The interior solution
Contrary to the exterior solution, the interior spacetime is not described by a vacuum solution. Here, for the sake of simplicity, we adopt the matter content inside a star given by the perfect fluid tensor
$T_{\mu\nu}^{M}=\left(\rho+p\right)u_{\mu}u_{\mu}+pg_{\mu\nu},$ (10)
where $\rho$ and $p$ are the energy density and pressure of the fluid, respectively, they are $r$-dependent functions. Also, the vector $u_{\mu}$ is the four-velocity of the fluid, which satisfies the normalization $u_{\mu}u^{\mu}=-1$. Above all, the interior solution described by a perfect fluid should match the vacuum exterior solution as $r=R$, where $R$ is the radius of the star. Due to the field (7), $B_{\mu\nu}=0$, and due to the matter content, one has $J_{B}^{\mu}=-J_{M}^{\mu}$ from Eq. (4). However, for the two cases studied here, we calculate that $|J_{B}^{\mu}|\approx 0$, consequently we assume that the matter sector is not coupled with the bumblebee field. Thus, the perfect fluid description (10) is still viable.
With the metric Ansatz (6) and the matter description (10), the modified field equations (3) are written as
$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $e^{2\beta}-(1+\ell)\left(1-2r\beta^{\prime}\right)\bigg{]}=\kappa\rho,$ (11)
$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $(1+\ell)\left(1+2r\alpha^{\prime}\right)-e^{2\beta}-\ell r^{2}\Big{(}\alpha^{\prime\prime}+\alpha^{\prime 2}$
$-\alpha^{\prime}\beta^{\prime}-\frac{2}{r}\beta^{\prime}\Big{)}\bigg{]}=\kappa p,$ (12)
$(1+\ell)e^{-2\beta}\bigg{[}$ $\alpha^{\prime\prime}+\alpha^{\prime 2}-\alpha^{\prime}\beta^{\prime}+\frac{1}{r}\left(\alpha^{\prime}-\beta^{\prime}\right)\bigg{]}=\kappa p.$ (13)
It is worth noticing that, due to $G_{\phi\phi}=\sin^{2}\theta G_{\theta\theta}$, one has just three independent equations from Eq. (3). Another equation comes from the local energy conservation equation $\nabla_{\nu}T^{\mu\nu}=0$. The component $\mu=r$ is not identically zero and will provide the modified TOV. First all, following the general relativity case, we assume that the metric component $g_{rr}$ for the exterior solution is
$g_{rr}=e^{2\beta}=(1+\ell)\left(1-\frac{2m(r)}{r}\right)^{-1},$ (14)
and the mass function in the spherically symmetric case is
$\frac{dm}{dr}=4\pi r^{2}\rho.$ (15)
In order to match the exterior solution, one assumes both $m(R)=M$ and $m(0)=0$. With the metric component (14), and with the aid of Eq. (13), the mentioned component of the energy-momentum equation conservation yields
$\frac{dp}{dr}=-\left(\frac{\rho+p+\ell\left(\rho+2p\right)}{1+2\ell}\right)\alpha^{\prime}-\frac{\ell\left(8\pi rp-m^{\prime\prime}\right)}{\left(1+2\ell\right)8\pi r^{2}}.$ (16)
This is the modified TOV equation in the bumblebee context for a star that matches the Schwarzschild-like solution for $r=R$, whose matter content is described by a perfect fluid. The TOV equation, whether the original or the modified ones, regards the hydrostatic equilibrium inside the star. A given solution of the TOV provides, from a specific equation of state, both the pressure and energy density configurations for a star.
From two components of the field equations, namely Eqs. (11) and (12), we obtain $\alpha^{\prime}$, which is given by
$\frac{d\alpha}{dr}=\frac{\left(1+2\ell\right)8\pi r^{3}p+\left(2+3\ell\right)m-\ell rm^{\prime}}{\left(2+3\ell\right)r(r-2m)}.$ (17)
Therefore, any interior metric should be calculated from Eqs. (16)-(17) by using a specific equation of state. In this work, as mentioned before, we focus on the incompressible case and quark stars with the MIT bag equation of state.
When the mass function $m(r)$ is previously specified, the (modified) TOV equation is an example of Riccati’s equation, a nonlinear differential equation. Such an equation could be solved from a linear second-order differential equation. [^1] See Ref. Ince , chapter 2, for more details. That is, if we write the modified TOV equation (16), with the aid of Eq. (17), as
$\frac{dp}{dr}=q_{0}+q_{1}p+q_{2}p^{2},$ (18)
it is possible to show that $p$ is solution of
$u^{\prime\prime}-q_{3}u^{\prime}+q_{4}u=0,$ (19)
in such a way that $q_{0},q_{1},q_{2},q_{3}=q_{1}+q^{\prime}_{2}/q_{2},$ and $q_{4}=q_{0}q_{2}$ are functions of the radial coordinate (except for $q_{2}$, their values will not be explicitly written here). The solution of (19)—in general large expressions in terms of hypergeometric functions—has two constants, which are fixed by the boundary conditions: (i) the pressure is zero on the star surface, $p(R)=0$, and (ii) $\alpha(R)=\frac{1}{2}\ln\left(1-\frac{2M}{R}\right)$ is necessary to match the exterior solution. From the trick of turning the nonlinear differential equation (16) into a linear and second-order equation (19), the pressure $p$ is simply
$p=-\frac{1}{q_{2}}\frac{u^{\prime}}{u}=\frac{(2+3\ell)(r-2m)}{8\pi r^{2}(1+2\ell)}\frac{u^{\prime}}{u}.$ (20)
As we can see, with a given $p$ and $m$, Eq. (17) provides the metric component $g_{tt}$. Assuming that $p$ comes from Eq. (20), the metric component is
[…]
It is worth emphasizing that the metric functions (14) and (21) are exact solutions of the modified field equations (11)-(13). In order to be an interior solution, $e^{2\beta}$ and $e^{2\alpha}$ should be positive definite and should keep the metric signature inside the star, i.e. for $r<R$. As we will see, that is the case for the two type of fluids presented here.
III Solving the modified TOV
III.1 Constant density star
The simplest case to be considered here is the incompressible fluid. In such a case, the energy density is constant inside the star, that is,
$\rho(r)=\rho_{0}.$ (22)
With a constant density, Eq (15) gives us $m(R)=M=\frac{4}{3}\pi R^{3}\rho_{0}$ for the total mass of the star. In the general relativity context, this simple case provides analytic expressions for the fluid pressure and the interior metric. This is the case in the bumblebee context, but solutions are much larger. As an alternative, we could make some approximations in order to get more reasonable expressions. For example, excluding terms like $\ell^{n}$ for $n>1$ and the second term on the right side of the modified TOV equation (16) (because $\ell\rho_{0}(p/\rho_{0}-1)\ll 1$), the function $u$, which is solution of Eq. (19), reads
$u(r)=\left(3-8\pi r^{2}\rho_{0}\right)^{\frac{2-3\ell}{8}}\left(C_{1}\left(3-8\pi r^{2}\rho_{0}\right)^{\frac{1}{2}}+C_{2}\right),$ (23)
where the two constants $C_{1}$ and $C_{2}$ are fixed from the boundary conditions using Eqs. (17) and (20), which connect $u$, $p$, and, consequently, $\alpha$. For the approximation adopted, one has
$C_{1}=$ $-\frac{\left(2-3\ell\right)}{4\sqrt{3}},$ (24)
$C_{2}=$ $\frac{\sqrt{3}}{4}(2-\ell)\left(3-8\pi R^{2}\rho_{0}\right)^{\frac{1}{2}}.$ (25)
The general relativity values are restored for $\ell=0$. Within this approximation, the fluid pressure is given by
$p(r)=\rho_{0}\Bigg{(}\frac{\sqrt{1-\frac{2M}{R}}-\sqrt{1-\frac{2Mr^{2}}{R^{3}}}}{\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}}\Bigg{)}$
$+\ell\rho_{0}\left(\frac{7-\frac{2M}{R}\left(\frac{r^{2}}{R^{2}}+6\right)-7\sqrt{\left(1-\frac{2Mr^{2}}{R^{3}}\right)\left(1-\frac{2M}{R}\right)}}{\left(\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}\right)^{2}}\right),$ (26)
which is the general relativity result plus a small contribution of the bumblebee field in this approximation. Notice also that the boundary condition $p(R)=0$ is satisfied. For $r=0$, the central pressure $p(0)=p_{c}$ diverges with $R=\frac{9}{4}M$, like the general relativity case. That is, any reasonable star should have $R>\frac{9}{4}M$ or, equivalently, $M<\frac{4R}{9}$ (this is the so-called Buchdahl’s limit in general relativity Buchdahl:1959zz ). Such a result is expected, for the bumblebee field does not change the event horizon radius in the exterior metric.
Fig. 1 shows the star pressure (exactly, numerically, and approximately calculated) with the same radius $R=3.2$ km (thus same mass) for different values of the Lorentz-violating parameter $\ell$. As we can see, the bumblebee field decreases the central pressure of the star. This will no longer be the case for the quark matter case in the next section. This effect in the constant-$\rho$ case is very counter-intuitive, because in this type of fluid the pressure decreases once the radius decreases. Using the approximate pressure (26) for $r=0$, it is possible to illustrate the relation between both the central pressure and the parameter $\ell$ in order to see the effect of the bumblebee field on $p_{c}$. In the same graphic (bottom), we see that the approximation (26) is no longer valid for $\ell>10^{-3}$, as we compare it with the pressure graphic (top).
Figure 1: Pressure (top) and central pressure (bottom) for a star with $\rho=\rho_{0}=2$ GeV/fm3 and $R=3.2$ km. In the pressure graphic, exact, numerical, and approximate calculations are compared each other. The bumblebee field decreases the central pressure of the star as $\ell$ increases. However, the approximation breaks for $\ell>10^{-3}$.
Figure 2: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using a constant density fluid. Exact, numerical, and approximate calculations are compared each other. We adopt $R=3.2$ km and $\rho=\rho_{0}=2$ GeV/fm3.
The interior metric (without any approximation) for constant energy density is straightforwardly calculated from Eq. (21) and $m(r)=\frac{4}{3}\pi r^{3}\rho_{0}$. That is,
$\alpha(r)=\ln\left[\frac{u(r)}{\left(3-8\pi r^{2}\rho_{0}\right)^{\frac{1}{4+6\ell}}}\right].$ (27)
Taking then the approximation for $u(r)$, given by Eq. (23), one has
$e^{\alpha(r)}=$ $\frac{3}{2}\sqrt{1-\frac{2M}{R}}-\frac{1}{2}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}+\frac{3\ell}{4}\bigg{(}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}$
$-\sqrt{1-\frac{2M}{R}}\bigg{)}.$ (28)
Even this approximation is in agreement with the boundary condition $\alpha(R)=\frac{1}{2}\ln\left(1-\frac{2M}{R}\right)$. As we can see in Fig. 2, for the exact, numerical, and approximate calculations, the metric components $e^{2\alpha}$ and $e^{2\beta}$ are positive definite inside the star. Thus, we have the correct metric signature both inside and outside the star.
Figure 3: Pressure (top), energy density (middle) and mass function (bottom) with the MIT bag equation of state. The bumblebee field, represented by the Lorentz-violating parameter $\ell$, increases the three quantities. We adopted $\mathcal{B}=60$ MeV/fm3 and $R=11$ km in these graphics.
Figure 4: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using the MIT bag model. We adopt $R=11$ km in these graphics.
III.2 Quark matter star
In the second type of fluid studied, we assume a star in the bumblebee gravity made up of quark matter with an equation of state more realistic than the one used in the constant-density star case. The quark matter could be described by the MIT bag model and by the following equation of state
$p=\frac{1}{3}\left(\rho-4\mathcal{B}\right),$ (29)
where, as we mentioned, $\mathcal{B}$ is the bag constant. The MIT bag model confines quarks inside a bag, and the bag’s energy, given as $E(r)=\frac{4\pi}{3}r^{3}\mathcal{B}+\frac{C}{r}$, combines a term proportional to the quarks’ kinetic energy and their volume. The bag energy takes place when the energy reaches its minimum, consequently determining the constant $C$. In this context, the bag constant is defined as $\mathcal{B}=E_{0}/(4V_{0})$, with $E_{0}$ representing the energy and $V_{0}$ denoting the volume of the bag. For a neutron, $m\approx E_{0}\approx 1\text{ GeV}$ with a radius of $r_{0}\approx 1\text{ fm}$, resulting in 60 $\mathcal{B}\approx 60\text{ GeV/fm}^{3}$ (or $8\times 10^{-5}$ km-2 in geometrized units) for numerical calculations. This value was used in Refs. Gondek-Rosinska:2008zmv ; Fraga:2013qra ; Arbanil:2016wud . However, given that the MIT bag model is a simple model, it is justifiable to slightly adjust this value. Thus, like Ref. Gondek-Rosinska:2008zmv , we also adopt $\mathcal{B}=40$ MeV/fm3 (equivalent to $5\times 10^{-5}$ km-2 in geometrized units). Unfortunately, for the equation of state (29) analytic solutions are not possible.
Once again, according to Fig. 3, like the constant density case, the bumblebee field modifies all physical parameters. Interestingly, contrary to the constant density fluid, given a star with same radius, the bumblebee model leads to higher central pressures when compared to the general relativity predictions. For example, for a star with $R=11.1$ km, using the MIT bag equation of state with $\mathcal{B}=60$ MeV/fm3, general relativity predicts $p_{c}=150$ Mev/fm3 for the central pressure. On the other hand, for $\ell=10^{-2}$, the bumblebee model gives us $p_{c}=160$ MeV/fm3. Consequently, the energy density also increases as the bumblebee field reaches the VEV. In this case, the explanation comes from the mass function $m(r)$, which also increases due to bumblebee field acting on the spacetime geometry (see Fig. 3 on the bottom).
Figure 5: Mass-radius relation in the bumblebee context for quark matter described by the MIT bag model with two values for the bag constant $\mathcal{B}$. For large values of the Lorentz-violating parameter $\ell$, the bumblebee gravity could provide larger mass-radius relations than general relativity. The pulsars PSR J0348+0432 Antoniadis:2013pzd $(M=2.01\pm 0.04M_{\odot})$, PSR J0740+6620 Fonseca:2021wxt $(M=2.08\pm 0.07M_{\odot})$, and PSR J0952-0607 Romani:2022jhd $(M=2.35\pm 0.17M_{\odot})$ are described/predicted by general relativity with the bag constant $\mathcal{B}=40$ MeV/fm3. On the other hand, the compact star in the GW 190814 LIGO event, with $M=2.59^{-0.08}_{-0.09}M_{\odot}$, is not. But it could be in the bumblebee context. See Table 1 for a more detailed comparison.
$\mathcal{B}=40$ MeV/fm3 $\ell=0$ $\ell=10^{-3}$ $\ell=1.9\times 10^{-2}$
$M(M_{\odot})$ 2.45 2.45 2.53
$R$ (km) 13.51 13.54 13.75
$p_{c}$ (MeV/fm3) 187.12 172.15 22.45
$\rho_{c}$ (MeV/fm3) 711.05 666.15 217.06
$\mathcal{B}=60$ MeV/fm3
$M(M_{\odot})$ 1.94 1.94 2.0
$R$ (km) 10.66 10.68 10.87
$p_{c}$ (MeV/fm3) 306.88 284.42 37.42
$\rho_{c}$ (MeV/fm3) 1,070.32 1,002.96 261.97
Table 1: Star parameters for values of the bag constant $\mathcal{B}$ and symmetry breaking parameter $\ell$. $M$ is the mass of the star, $R$ is its radius, $p_{c}$ is the central pressure, and $\rho_{c}$ is the central density.
Like the constant-density case, Fig. 4 shows the metric functions $e^{\alpha}$ and $e^{\beta}$ are positive definite inside the star $(r<R)$, consequently the metric is well behaved and its signature (,+,+,+)(-,+,+,+)( - , + , + , + ) is also preserved.
As mentioned in Introduction, the bumblebee model could provide more massive stars than the general relativity context, for the mass-radius relation is also modified due to the bumblebee field and its VEV. As we can see in Fig. 5, large values for the Lorentz breaking parameter $\ell$ increases the mass-radius relation dramatically, except for a threshold (about $\ell=10^{-1}$) from which larger values of $\ell$ do not increase that relation. Therefore, a modified gravitational model like the bumblebee model could fit data for large mass-radius relations in which the general relativity context is not able to fit. In particular, we see in Fig. 5 and Table 1 that for $\mathcal{B}=60$ MeV/fm3 and $\ell=1.9\times 10^{-2}$, the bumblebee model give us a star with $R=2M_{\odot}$. On the other hand, general relativity is not able to reach the $R=2M_{\odot}$ limit. With a small bag constant (as we can also see in Table 1), for example $\mathcal{B}=40$ MeV/fm3, the bumblebee model is able to describe even more massive stars, like the one in the gravitational wave event GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$. Contrary to general relativity, which does not fit this data ($M=2.45M_{\odot}$), the bumblebee model with $\ell=1.9\times 10^{-2}$ predicts a massive quark star with $M=2.53M_{\odot}$ and $R=13.75$ km.
IV Final remarks
The bumblebee gravity is a Lorentz-violating model in which the VEV of the bumblebee field is nonzero, modifying then fields and spacetime. Within this context of modified theory of gravity, spherically symmetry metrics, describing the interior of noncharged and nonrotating stars, were obtained analytically.
By using two types of fluids, we show that the bumblebee field could increase or decrease the fluid pressure inside the star. In particular for the more realistic fluid studied in this article, quark matter described by the MIT bag model, the Lorentz violation increases both the pressure and the energy density of the fluid of quark stars with same radius.
In the end of the day, it is reasonable to hope for the Lorentz breaking symmetry in contexts of high energy and intense gravitational effects. The results presented in this article point out to larger mass-radius relations in the bumblebee context when compared to the general relativity case. In this sense, recent data of large mass-radius relations, like the gravitational wave events, GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$, could be justified by a modified gravity model like the bumblebee model using quark matter as the star matter content.
Recently, three works on bumblebee gravity and stars drawn our attention. The first one Ji:2024aeg builds an interior solution from another exterior solution in bumblebee gravity, namely the spacetime obtained in Bertolami and Páramos Bertolami . In our case, as mentioned before, the exterior solution is the Schwarzschild-like metric of Casana et al. Casana . In this case, there is no overlapping results. Regarding the two other papers, those authors tried to obtain the interior solution of a star from the Schwarzschild-like metric. The modified TOV is incorrect in either cases. In Ref. Liu , the polytropic fluid studied is not solution of all modified field equations, and in Ref. Panotopoulos:2024jtn even the exterior solution is not solution of the field equations in the bumblebee context.
Acknowledgments
JCSN thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, (Grant No. 170579/2023-9) for the financial support and the ICT-Unifal for the kind hospitality. F.G.G. is supported by CNPq through 306762/2021-8.

### v3

@@
 When compared to the general relativity mass-radius relation, a Lorentz symmetry breaking context, like the bumblebee gravity, could provide more massive stars, surpassing the $2.5M_{\odot}$ limit as the interior of the star is described by quark matter with the MIT bag model.
+Also, we investigate the stability of the solution with the MIT bag equation of state in this context of modified gravity.
 Neutron Star, Quark Gluon Plasma, Lorentz Symmetry Breaking
@@
 In this sense, quark matter in astrophysics has been subject of many papers.
-In modified theories of gravity, we see works in Einstein-Gauss-Bonnet gravity Gammon:2023uss , $f(R,L_{m},T)$ gravity Tangphati:2024war , Weyl geometric gravity Haghani:2023nrm and many others (see Ref.
+In modified theories of gravity, we see works in Einstein-Gauss-Bonnet gravity Tangphati:2021tcy ; Tangphati:2021mvu ; Pretel:2021czp ; Gammon:2023uss ; Pretel:2021gqq , $f(R,L_{m},T)$ gravity Tangphati:2024war , Weyl geometric gravity Haghani:2023nrm and many others (see Ref.
 Olmo:2019flu for a review).
@@
 On the other hand, a star in the bumblebee gravity could have $M=2.53M_{\odot}$ or even larger values.
+It is worth pointing out that the nature of the compact object in the GW190814 event is still unknown.
+Besides the quark star hypothesis Tangphati:2024war ; Haghani:2023nrm , there are alternatives that describe such a compact object as a light black hole LIGO ; Nathanail:2021tay , a fast rotating neutron star Zhang:2020zsc ; Most:2020bba ; Biswas:2020xna or even a hybrid star Zhang:2020dfi .
+Another alternative is to describe the compact object in the GW190814 event as a neutron star in the modified context of the Eddington-inspired Born-Infeld theory Prasetyo:2021kfx .
+In such a context, a neutron star with $2.6M_{\odot}$ is possible.
+However, as mentioned, we assume here that the compact object in the GW190814 event could be a quark star in bumblebee gravity.
+Recently, three works on bumblebee gravity and stars drawn our attention.
+The first one Ji:2024aeg builds an interior solution from another exterior solution in bumblebee gravity, namely the spacetime obtained in Bertolami and Páramos Bertolami .
+In our case, as mentioned before, the exterior solution is the Schwarzschild-like metric of Casana et al.
+Casana .
+In this case, due to the fact that different exterior solutions generate different interior solutions, there are no overlapping results.
+Regarding the two other papers, the authors tried to obtain the interior solution of a star from the Schwarzschild-like metric like us.
+But the modified TOV is incorrect in either cases.
+In Ref.
+Liu , for example, the polytropic fluid studied is not solution of all modified field equations, and in Ref.
+Panotopoulos:2024jtn even the exterior solution (a vacuum solution) is not solution of the field equations obtained in the mentioned work.
 This article is structured as follows: In Section II the adopted bumblebee model is presented, the exterior solution, Schwarzschild-like spacetime, is briefly discussed.
@@
 The mass-radius result is exhibited, and we can see the enhancement of this relation due to the bumblebee field.
-The final comments are in Section IV.
+In Section IV, stability of the solution with the MIT bag equation of state is studied calculating the speed of sound in the fluid, the adiabatic index and the mass-central pressure diagram.
+The final comments are in Section V.
 In our calculations we adopt geometrized units, i.e., $G=c=1$, where $G$ is the Newtonian constant, and $c$ is speed of light in vacuum.
@@
 The model adopted here regards a vector field $B_{\mu}$, the bumblebee field (in a torsion-free spacetime), given by the following action:
-$S_{B}$ $=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2\kappa}+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}$
-$\ \ \ -V(B^{\mu}B_{\mu}\pm b^{2})+\mathcal{L}_{M}\bigg{]},$ (1)
+$S_{B}=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2\kappa}+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})+\mathcal{L}_{M}\bigg{]},$ (1)
 where $\kappa=8\pi G/c^{4}$ is the gravitational coupling constant, and $\xi$ (with mass dimension $[\xi]=M^{-1}$) plays the role of a coupling constant that accounts for the nonminimum interaction between the bumblebee field and the geometry Bailey2006 .
@@
 $G_{\mu\nu}=$ $\ R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa\left(T^{B}_{\mu\nu}+T^{M}_{\mu\nu}\right)$
-$=$ $\ \kappa\left[2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu}\right]$
-$+\xi\bigg{[}\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}$
-$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
-$-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\bigg{]}+\kappa T^{M}_{\mu\nu}.$ (3)
+$=$ $\ \kappa\left[2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu}\right]+\xi\bigg{[}\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}$
+$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\bigg{]}+\kappa T^{M}_{\mu\nu},$ (3)
 The tensor $G_{\mu\nu}$ is the usual Einstein tensor, and the operator means derivative with respect to the potential argument (consequently, throughout this article is also derivative with respect to the radial coordinate), and $\nabla_{\mu}$ is the covariant derivative.
@@
 Casana , the metric able to describe a static and neutral black hole in vacuum (or alternatively the exterior region of a nonrotating star) in the bumblebee context is
-$ds^{2}=$ $-\left(1-\frac{2M}{r}\right)dt^{2}+(1+\ell)\left(1-\frac{2M}{r}\right)^{-1}dr^{2}$
-$+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (9)
+$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+(1+\ell)\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (9)
 That is to say, the spacetime described by Eq.
@@
 $\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $e^{2\beta}-(1+\ell)\left(1-2r\beta^{\prime}\right)\bigg{]}=\kappa\rho,$ (11)
-$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $(1+\ell)\left(1+2r\alpha^{\prime}\right)-e^{2\beta}-\ell r^{2}\Big{(}\alpha^{\prime\prime}+\alpha^{\prime 2}$
-$-\alpha^{\prime}\beta^{\prime}-\frac{2}{r}\beta^{\prime}\Big{)}\bigg{]}=\kappa p,$ (12)
+$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $(1+\ell)\left(1+2r\alpha^{\prime}\right)-e^{2\beta}-\ell r^{2}\Big{(}\alpha^{\prime\prime}+\alpha^{\prime 2}-\alpha^{\prime}\beta^{\prime}-\frac{2}{r}\beta^{\prime}\Big{)}\bigg{]}=\kappa p,$ (12)
 $(1+\ell)e^{-2\beta}\bigg{[}$ $\alpha^{\prime\prime}+\alpha^{\prime 2}-\alpha^{\prime}\beta^{\prime}+\frac{1}{r}\left(\alpha^{\prime}-\beta^{\prime}\right)\bigg{]}=\kappa p.$ (13)
@@
 The general relativity values are restored for $\ell=0$. Within this approximation, the fluid pressure is given by
-$p(r)=\rho_{0}\Bigg{(}\frac{\sqrt{1-\frac{2M}{R}}-\sqrt{1-\frac{2Mr^{2}}{R^{3}}}}{\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}}\Bigg{)}$
-$+\ell\rho_{0}\left(\frac{7-\frac{2M}{R}\left(\frac{r^{2}}{R^{2}}+6\right)-7\sqrt{\left(1-\frac{2Mr^{2}}{R^{3}}\right)\left(1-\frac{2M}{R}\right)}}{\left(\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}\right)^{2}}\right),$ (26)
+$p(r)=\rho_{0}\left(\frac{\sqrt{1-\frac{2M}{R}}-\sqrt{1-\frac{2Mr^{2}}{R^{3}}}}{\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}}\right)+\ell\rho_{0}\left(\frac{7-\frac{2M}{R}\left(\frac{r^{2}}{R^{2}}+6\right)-7\sqrt{\left(1-\frac{2Mr^{2}}{R^{3}}\right)\left(1-\frac{2M}{R}\right)}}{\left(\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}\right)^{2}}\right),$ (26)
 which is the general relativity result plus a small contribution of the bumblebee field in this approximation.
@@
 Such a result is expected, for the bumblebee field does not change the event horizon radius in the exterior metric.
+Figure 1: Pressure (left) and central pressure (right) for a star with $\rho=\rho_{0}=2$ GeV/fm3 and $R=3.2$ km.
+In the pressure graphic, exact, numerical, and approximate calculations are compared each other.
+The bumblebee field decreases the central pressure of the star as $\ell$ increases.
+However, the approximation breaks for $\ell>10^{-3}$.
+Figure 2: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using a constant density fluid.
+Exact, numerical, and approximate calculations are compared each other.
+We adopt $R=3.2$ km and $\rho=\rho_{0}=2$ GeV/fm3.
 Fig.
@@
 Using the approximate pressure (26) for $r=0$, it is possible to illustrate the relation between both the central pressure and the parameter $\ell$ in order to see the effect of the bumblebee field on $p_{c}$.
-In the same graphic (bottom), we see that the approximation (26) is no longer valid for $\ell>10^{-3}$, as we compare it with the pressure graphic (top).
-Figure 1: Pressure (top) and central pressure (bottom) for a star with $\rho=\rho_{0}=2$ GeV/fm3 and $R=3.2$ km.
-In the pressure graphic, exact, numerical, and approximate calculations are compared each other.
-The bumblebee field decreases the central pressure of the star as $\ell$ increases.
-However, the approximation breaks for $\ell>10^{-3}$.
-Figure 2: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using a constant density fluid.
-Exact, numerical, and approximate calculations are compared each other.
-We adopt $R=3.2$ km and $\rho=\rho_{0}=2$ GeV/fm3.
+In the same graphic (right), we see that the approximation (26) is no longer valid for $\ell>10^{-3}$, as we compare it with the pressure graphic (left).
 The interior metric (without any approximation) for constant energy density is straightforwardly calculated from Eq. (21) and $m(r)=\frac{4}{3}\pi r^{3}\rho_{0}$. That is,
@@
 Taking then the approximation for $u(r)$, given by Eq. (23), one has
-$e^{\alpha(r)}=$ $\frac{3}{2}\sqrt{1-\frac{2M}{R}}-\frac{1}{2}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}+\frac{3\ell}{4}\bigg{(}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}$
-$-\sqrt{1-\frac{2M}{R}}\bigg{)}.$ (28)
+$e^{\alpha(r)}=\frac{3}{2}\sqrt{1-\frac{2M}{R}}-\frac{1}{2}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}+\frac{3\ell}{4}\bigg{(}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-\sqrt{1-\frac{2M}{R}}\bigg{)}.$ (28)
 Even this approximation is in agreement with the boundary condition $\alpha(R)=\frac{1}{2}\ln\left(1-\frac{2M}{R}\right)$.
@@
 Thus, we have the correct metric signature both inside and outside the star.
-Figure 3: Pressure (top), energy density (middle) and mass function (bottom) with the MIT bag equation of state.
-The bumblebee field, represented by the Lorentz-violating parameter $\ell$, increases the three quantities.
-We adopted $\mathcal{B}=60$ MeV/fm3 and $R=11$ km in these graphics.
-Figure 4: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using the MIT bag model.
-We adopt $R=11$ km in these graphics.
 III.2 Quark matter star
@@
 In this context, the bag constant is defined as $\mathcal{B}=E_{0}/(4V_{0})$, with $E_{0}$ representing the energy and $V_{0}$ denoting the volume of the bag.
-For a neutron, $m\approx E_{0}\approx 1\text{ GeV}$ with a radius of $r_{0}\approx 1\text{ fm}$, resulting in 60 $\mathcal{B}\approx 60\text{ GeV/fm}^{3}$ (or $8\times 10^{-5}$ km-2 in geometrized units) for numerical calculations.
-This value was used in Refs.
+For a neutron, $m\approx E_{0}\approx 1\text{ GeV}$ with a radius of $r_{0}\approx 1\text{ fm}$, resulting in 60 $\mathcal{B}\approx 60\text{ MeV/fm}^{3}$ (or $8\times 10^{-5}$ km-2 in geometrized units) for numerical calculations.
+This value falls within the range of the bag constant, $\mathcal{B}\sim 57-92$ MeV/fm3, allowing for the reproduction of hadronic masses at low-energy regimes DeTar:1983rw , and has been utilized in Refs.
 Gondek-Rosinska:2008zmv ; Fraga:2013qra ; Arbanil:2016wud .
-However, given that the MIT bag model is a simple model, it is justifiable to slightly adjust this value.
+However, given that the MIT bag model is a simple model, in extreme energy regimes the strong coupling constant for quark interactions is less than that one utilized in DeTar:1983rw .
+In extreme density regimes, the strong coupling constant is less than in typical conditions where values of $\mathcal{B}$ were derived.
+In this sense, it is justifiable to slightly adjust the bag constant.
 Thus, like Ref.
@@
 Unfortunately, for the equation of state (29) analytic solutions are not possible.
+Figure 3: Pressure (left), energy density (right) and mass function (bottom) with the MIT bag equation of state.
+The bumblebee field, represented by the Lorentz-violating parameter $\ell$, increases the three quantities.
+We adopted $\mathcal{B}=60$ MeV/fm3 and $R=11$ km in these graphics.
+Figure 4: Positivity of the metric functions, $g_{tt}=e^{2\alpha}$ and $g_{rr}=e^{2\beta}$, for very exaggerated values of the Lorentz-violating parameter $\ell$ using the MIT bag model.
+We adopt $R=11$ km in these graphics.
 Once again, according to Fig.
@@
 3 on the bottom).
-Figure 5: Mass-radius relation in the bumblebee context for quark matter described by the MIT bag model with two values for the bag constant $\mathcal{B}$.
+Like the constant-density case, Fig.
+4 shows the metric functions $e^{\alpha}$ and $e^{\beta}$ are positive definite inside the star $(r<R)$, consequently the metric is well behaved and its signature (,+,+,+)(-,+,+,+)( - , + , + , + ) is also preserved.
+As mentioned in Introduction, the bumblebee model could provide more massive stars than the general relativity context, for the mass-radius relation is also modified due to the bumblebee field and its VEV.
+As we can see in Fig.
+5, large values for the Lorentz-violating parameter $\ell$ increases the mass-radius relation dramatically, except for a threshold (about $\ell=10^{-1}$) from which larger values of $\ell$ do not increase that relation.
+Also, it is worth mentioning the “incomplete” curves for large values of $\ell$ in Fig.
+5.
+A given curve in the mass-radius graphic is solution of the TOV with conditions like $p(r)\geq 0$ and $r-2m(r)\geq 0$.
+For large values of the Lorentz-violating parameter, those conditions are satisfied just for smaller range of parameters.
+As we will see, this is linked to the stability of the solutions.
+Therefore, for appropriate values of the Lorentz-violating parameter $\ell$, a modified gravitational model like the bumblebee model could fit data for large mass-radius relations in which the general relativity context is not able to fit.
+In particular, we see in Fig.
+5 and Table 1 that for $\mathcal{B}=60$ MeV/fm3 and $\ell=1.9\times 10^{-2}$, the bumblebee model give us a star with $R=2M_{\odot}$.
+On the other hand, general relativity is not able to reach the $R=2M_{\odot}$ limit.
+With a small bag constant (as we can also see in Table 1), for example $\mathcal{B}=40$ MeV/fm3, the bumblebee model is able to describe even more massive stars, like the one in the gravitational wave event GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$.
+Contrary to general relativity, which does not fit this data ($M=2.45M_{\odot}$), the bumblebee model with $\ell=1.9\times 10^{-2}$ predicts a massive quark star with $M=2.53M_{\odot}$ and $R=13.75$ km.
+This is a virtue of modified theories of gravity Gammon:2023uss ; Tangphati:2024war ; Haghani:2023nrm .
+Because theories that predict smaller mass-radius relations than general relativity does would be under pressure due to recent data.
+Figure 5: On the left, the mass-radius relation in the bumblebee context for quark matter described by the MIT bag model with two values for the bag constant $\mathcal{B}$.
 For large values of the Lorentz-violating parameter $\ell$, the bumblebee gravity could provide larger mass-radius relations than general relativity.
@@
 See Table 1 for a more detailed comparison.
+On the right, the $M-\rho_{c}$ curve shows stability just for $\frac{dM}{d\rho_{c}}>0$.
+The black dots $(\frac{dM}{d\rho_{c}}=0)$ indicate the limit of stable configurations.
 $\mathcal{B}=40$ MeV/fm3 $\ell=0$ $\ell=10^{-3}$ $\ell=1.9\times 10^{-2}$
@@
 $M$ is the mass of the star, $R$ is its radius, $p_{c}$ is the central pressure, and $\rho_{c}$ is the central density.
-Like the constant-density case, Fig.
-4 shows the metric functions $e^{\alpha}$ and $e^{\beta}$ are positive definite inside the star $(r<R)$, consequently the metric is well behaved and its signature (,+,+,+)(-,+,+,+)( - , + , + , + ) is also preserved.
-As mentioned in Introduction, the bumblebee model could provide more massive stars than the general relativity context, for the mass-radius relation is also modified due to the bumblebee field and its VEV.
+IV Stability analysis
+In this section, we briefly comment stability of the bumblebee star with quark matter described by the MIT bag equation of state (29).
+In general relativity, a necessary but not sufficient stability criterion regards the central pressure $\rho_{c}$.
+It is assumed, even in modified gravity works, that $\frac{dM}{d\rho_{c}}>0$ is a necessary condition to stability in the static case Zeldovich ; Olmo:2019flu .
 As we can see in Fig.
-5, large values for the Lorentz breaking parameter $\ell$ increases the mass-radius relation dramatically, except for a threshold (about $\ell=10^{-1}$) from which larger values of $\ell$ do not increase that relation.
-Therefore, a modified gravitational model like the bumblebee model could fit data for large mass-radius relations in which the general relativity context is not able to fit.
-In particular, we see in Fig.
-5 and Table 1 that for $\mathcal{B}=60$ MeV/fm3 and $\ell=1.9\times 10^{-2}$, the bumblebee model give us a star with $R=2M_{\odot}$.
-On the other hand, general relativity is not able to reach the $R=2M_{\odot}$ limit.
-With a small bag constant (as we can also see in Table 1), for example $\mathcal{B}=40$ MeV/fm3, the bumblebee model is able to describe even more massive stars, like the one in the gravitational wave event GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$.
-Contrary to general relativity, which does not fit this data ($M=2.45M_{\odot}$), the bumblebee model with $\ell=1.9\times 10^{-2}$ predicts a massive quark star with $M=2.53M_{\odot}$ and $R=13.75$ km.
-IV Final remarks
+5, the $M-\rho_{c}$ curve indicates stability intervals for both different values of the bag constant $\mathcal{B}$ and the Lorentz-violating parameter $\ell$.
+Also we can see that large values of $\ell$ yield small stable ranges.
+This is in agreement with the mass-radius diagram that shows “incomplete” curves for large values of $\ell$.
+It is common to adopt the speed of sound in the fluid, $c_{s}=\sqrt{\frac{dp}{d\rho}}$, to check whether or not causality and stability issues could arise.
+Inside a stable star, the speed of sound of the fluid must never surpass the speed of light $c$.
+Using the MIT bag equation of state, it is easy to see that $c_{s}=\frac{c}{\sqrt{3}}$ (where the speed of light in vacuum was restored), that is, one has a constant subluminal speed of sound in the fluid, thus no causality and stability issues are present due to the speed of sound.
+Another useful parameter to investigate the star stability is the adiabatic index of perturbations, namely
+$\gamma=\left(1+\frac{\rho}{p}\right)\left(\frac{dp}{d\rho}\right)_{S},$ (30)
+where $S$ means that the derivative (or the speed of sound in the fluid) is calculated at a constant specific entropy.
+In the general relativity context, for spherical symmetry, Chandrasekhar Chandrasekhar:1964zz showed using radial perturbations that $\gamma_{cr}=\frac{4}{3}+\frac{19}{42}b$ is the critical value of the adiabatic index for a given stable star, where $b$ is the compactness parameter.
+That is, stable stars must have $\gamma>\gamma_{cr}$.
+In bumblebee gravity (and other modified gravity theories) there is no available $\gamma_{cr}$ yet.
+Thus, it is common to assume that stable stars in these alternative contexts must satisfy at least $\gamma>\frac{4}{3}$, which is the Newtonian limit for the adiabatic index.
+This is the case here, for using the MIT bag equation of state (29), one has $\gamma=\frac{1}{3}\left(4+\frac{1}{p}\right)>\frac{4}{3}$, according to the positivity of $p$ in Fig.
+3, suggesting then stability.
+V Final remarks
 The bumblebee gravity is a Lorentz-violating model in which the VEV of the bumblebee field is nonzero, modifying then fields and spacetime.
@@
 In this sense, recent data of large mass-radius relations, like the gravitational wave events, GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$, could be justified by a modified gravity model like the bumblebee model using quark matter as the star matter content.
-Recently, three works on bumblebee gravity and stars drawn our attention.
-The first one Ji:2024aeg builds an interior solution from another exterior solution in bumblebee gravity, namely the spacetime obtained in Bertolami and Páramos Bertolami .
-In our case, as mentioned before, the exterior solution is the Schwarzschild-like metric of Casana et al.
-Casana .
-In this case, there is no overlapping results.
-Regarding the two other papers, those authors tried to obtain the interior solution of a star from the Schwarzschild-like metric.
-The modified TOV is incorrect in either cases.
-In Ref.
-Liu , the polytropic fluid studied is not solution of all modified field equations, and in Ref.
-Panotopoulos:2024jtn even the exterior solution is not solution of the field equations in the bumblebee context.
+Another interesting point regards the stability analysis.
+The $M-\rho_{c}$ curve—from the $\frac{dM}{d\rho_{c}}>0$ stability criterion—indicates stability for the bumblebee stars with quark matter.
+Moreover, large values of the Lorentz-violating parameter $\ell$ provide larger mass-radius relations, but the range of allowed and stable configurations decreases with that parameter.
 Acknowledgments
@@
 is supported by CNPq through 306762/2021-8.
+The authors also thank two anonymous referees for valuable comments.